OpenStudy (anonymous):

9/5+ (-4/3) how do I solve?

OpenStudy (anonymous):

since you have common denominators all you need to do now is subtract the nominators 1-5 which will give you -4 So your answer will be \[\frac{ -4 }{ 3 }\]

OpenStudy (anonymous):

thks need help with 3 others (1/3)+ 3/8?

OpenStudy (anonymous):

you need common denominators before you add these so you find a number that both 3 and 8 go into and we find 24. Three goes into 24 eight times so we multiply both the bottom and top by 8. Do the same for 8 except 8 goes into 24 three times so we multiply both the top and bottom by 3. So we get \[\frac{ 8 }{ 24 } + \frac{ 9 }{ 24 }\] which gives us \[\frac{ 17 }{ 24 }\]

OpenStudy (anonymous):

don't forget to choose a best response :)

OpenStudy (anonymous):

thks. and how about (-10/7) + 1/6

OpenStudy (anonymous):

It's the exact same process, except now you have to go a lot higher to find a number that both 7 and 6 go into and you will eventually find 42 \[\frac{ -60 }{ 42 } +\frac{ 7 }{ 42 }\] add -60 and 7 to get \[\frac{ -53 }{ 42 }\]

OpenStudy (anonymous):

how to solve 9/5+ (- 4/3)

OpenStudy (anonymous):

@max0236 Everything seems very good. However, unless the original question was deleted, on the first question, the fractions do not have common denominators. :/